OpenStudy (anonymous):

Can you tell me if my answer is correct? I am being asked to flip a coin 4 times and determine what the probability of each of my coin flips turning up tails or heads is. This is my answer. The probability of each of my coin flips turning up “male” or “female” is C (4, 2) = 4! / (4 - 2)! (2!) = 4*3*2*1 / (2*1) (2*1) = 24 / 2*2 = 24/4 = 6 Total number of possibilities = 2^n where n = number of events (flips) So, total = 2^4 = 16 So probability = (number of times two tails appear) / (total number of events) = 6/16 = 3/8 Did I solve that correctly?

Directrix (directrix):

Take a look at this video after you get this problem solved. Good reinforcement. http://www.youtube.com/watch?v=xNLQuuvE9ug Here is work based on my understanding of the problem: We can get 0T, 1T, 2T, 3T, or 4T. P(0T) = C(4,0) (1/2)^0 ((1/2)^4 = 1 x 1 x (1/16) = 1/16 wher C(4,0) means a set of 4 taken 0 at the time or four choose zero. P(1T) = C(4,1) (1/2)^1 (1/2)^3 = 4 x 1/2 x 1/8 = 4/16 P(2T) = C(4,2) (1/2)^2(1/2)^2 = 6 x1/4 x 1/4 = 6/16 P(3T) = C(4,3) (1/2)^3 (1/2)^1 = 4 x 1/8 x 1/2 = 4/16 P(4T) = C4,4) (1/2)^4 (1/2)^0 = 1 x 1/16 x 1 = 1/16 Note that you are most likely to get two tails but that is not certain.